# Unlocking the Potential of Model Merging for Low-Resource Languages

### v2

@@ Abstract
Adapting large language models (LLMs) to new languages typically involves continual pre-training (CT) followed by supervised fine-tuning (SFT). However, this CT-then-SFT approach struggles with limited data in the context of low-resource languages, failing to balance language modeling and task-solving capabilities. We thus propose model merging as an alternative for low-resource languages, combining models with distinct capabilities into a single model without additional training. We use model merging to develop task-solving LLMs for low-resource languages without SFT data in the target languages. Our experiments based on Llama-2-7B demonstrate that model merging effectively endows LLMs for low-resource languages with task-solving abilities, outperforming CT-then-SFT in scenarios with extremely scarce data. Observing performance saturation in model merging with more training tokens, we further analyze the merging process and introduce a slack variable to the model merging algorithm to mitigate the loss of important parameters, thereby enhancing performance. We hope that model merging can benefit more human languages suffering from data scarcity with its higher data efficiency.
@@ 1 Introduction
To answer these questions, we study the adaptation of Llama-2-7B Touvron et al. (2023), an English-centric LLM, into seven distinct low-resource languages. We first continually pre-train Llama-2-7B on monolingual texts in each language. Next, we explore two approaches to inject task-solving capabilities into this continually pre-trained model: (1) training the LLM with English SFT data or the data translated to the target low-resource language; (2) merging the model with an English task-solving LLM. Experiments show that model merging can effectively equip the CT models with task-solving capabilities. Notably, when pre-training corpora in the target language are extremely scarce (<10B tokens), model merging outperforms CT-then-SFT. As an LLM is pre-trained with more tokens, the improvements brought by model merging gradually saturate. Then, model merging can no longer significantly surpass the SFT method.
To further investigate the factors impeding the continuous improvement of model merging, we conduct a detailed analysis of the process of merging two LLMs. We find that when an English SFT model is merged with an LLM continually pretrained with more tokens in the target language, more parameters from the SFT model are discarded during merging. The loss of these parameters may lead to a decline in task-solving capabilities, preventing the merged model from improving performance on downstream tasks. To mitigate the loss of important parameters from the SFT model, we propose model merging with slack variables. This strategy allows for more flexible control over the merging process to retain the important parameters.
@@ 3 Model Merging for Low-Resource Languages
The conventional CT-then-SFT paradigm struggles to balance language modeling and task-solving abilities in the context of low-resource languages. We propose model merging as an alternative, which can construct task-solving LLMs for low-resource languages without requiring SFT data in the target languages.
@@ 3.2 Roadmap Towards LLMs for Low-Resource Languages
Given a base model only pre-trained on an English-centric corpus, e.g., Llama-2-7B Touvron et al. (2023) in our study, we want to construct a model capable of solving tasks in a low-resource language. For the target language, there are very limited pre-training texts, ranging from 1B to 20B tokens, and almost no data for supervised finetuning (SFT). In this scenario, we investigate two representative paradigms of constructing such a model: CT-then-SFT and model merging. We illustrate the roadmap in Figure 1, which demonstrates the relations between the models.
@@
The common practice is (1) first continual pre-training (CT) on the monolingual texts in the target language X to learn the language modeling and (2) then learning task-solving abilities through SFT Yong et al. (2023); Nguyen et al. (2023). This approach is referred to as CT-then-SFT. In this approach, we have the following models:
@@
(2) CTSFT-X-mt: We translate FLAN and the training set of GSM8K into the target language X with machine translation (MT) systems [^2] We use NLLB NLLB Team et al. (2022) for translation. To enhance the model’s ability to follow English prompts, we randomly sample half of the training instances for translation. The other half of instances remain in English., which is a common practice to obtain SFT data for non-English languages Muennighoff et al. (2023); Li et al. (2023a, b). We then finetune CT-X with the translated data.
@@
By model merging, we can integrate distinct LLMs with various capabilities into one LLM. To obtain a model capable of solving task in the target language X, we can merge the following two models:
CT-X: As discussed in the CT-then-SFT procedure, this model learned a certain amount of language modeling in the language X after CT. However, its task-solving ability is limited.
@@ 4 Experimental Setup
We select these languages because they are underrepresented in currently popular LLMs despite their large population (over 475M) worldwide. As shown in Table 3, the performance of Llama-2-7B in these languages is close to or even worse than random guessing. Notably, the vocabulary of Llama-2 does not even contain tokens for Odia and traditional Mongolian, which indicates that the model has hardly seen these languages during pre-training. Moreover, limited resources are available for these languages on the internet. Among those languages, we can only collect fewer than 1B tokens of monolingual texts for Odia, Tibetan, Uyghur, and traditional Mongolian. The problem of data scarcity becomes more severe in terms of high-quality data for supervised fine-tuning.
@@
We translate FLAN into the languages of our study using NLLB-200-Distilled-1.3B NLLB Team et al. (2022) [^3] NLLB is currently the open-source MT model with the most extensive language support.. Note that this model does not support traditional Mongolian and there are no open-source MT models available for this language currently. We thus adopt a roundabout way: translating the instructions into Cyrillic Mongolian, which NLLB supports, and converting them into traditional Mongolian with an open-source transliteration tool [^4] https://github.com/tugstugi/mongolian-nlp/tree/master/bichig2cyrillic.
@@
Table 2: The pertaining corpora and evaluation tasks of each language in our study. The number of tokens is obtained with the tokenizer of Llama-2.
@@
Table 3: Performance of models built through the roadmap. The best results are made bold, with the second underlined. The Task and Lang. columns denotes whether the model have enhanced task-solving abilities and learned the target languages, respectively.
@@ 5.1 Effectiveness of Model Merging
In the conventional CT-then-SFT approach, LLMs often fail to acquire sufficient comprehension of the target language from a small-scale CT corpus, and this ability may be further diminished by supervised fine-tuning. In contrast, model merging can preserve the language modeling acquired during CT while incorporating task-solving capabilities by resolving parameter conflicts carefully.
@@ 5.2 Performance Plateau of Model Merging
Although model merging proves more effective than CT-then-SFT in most languages of our study, this is not the case for Bengali, which utilizes the largest corpus in the experiments: CTSFT-ben-flan outperforms TIES-ben-flan. We guess that the applicability of model merging may be related to the amount of corpus used in CT. Thus, we examine the performance changes of the merged models and the CT-then-SFT models under different amounts of CT tokens.
We collect the intermediate checkpoints of CT-ben and CT-tel, which are the two languages with the largest size of pre-training corpora. Then, we derive CTSFT-X-flan and TIES-X-flan models based on these checkpoints.
Figure 2 illustrates the performance of CT-only (CT-X), CT-then-SFT (CTSFT-X-flan), and model merging (TIES-X-flan) based on every checkpoint. For Bengali and Telugu, TIES-X-flan outperforms CT-X in each checkpoint, indicating model merging can robustly enhance LLM’s task-solving ability, while CT-then-SFT demonstrates greater variability. And in both languages, for all checkpoints where the amount of pre-trained tokens is less than 10.4B tokens, TIES-X-flan can achieve better results than CTSFT-X-flan.
Model merging can integrate language modeling and task-solving capabilities more effectively than CT-then-SFT, in scenarios with limited language resources ($<$10B tokens). Four out of the seven languages in our study utilize corpora with fewer than 10B tokens. This data scarcity is characteristic of low-resource languages. For instance, 81% (135 out of 166) of the languages in the multilingual corpus CulturaX Nguyen et al. (2024) have fewer than 10B tokens. Consequently, our findings on the effectiveness of model merging are broadly applicable and have the potential to benefit a wide range of human languages.
As the amount of CT tokens increases, the CTSFT-X-flan models show more rapid improvement in task-solving capabilities compared to TIE-X-flan. Specifically, when pre-trained with more than 14.4B tokens, CTSFT-ben-flan demonstrates superior performance over TIES-ben-flan. Similarly, in Telugu, the performance gap between the two models diminishes with additional CT tokens. However, due to the smaller size of the Telugu corpus compared to Bengali’s, we have not observed CTSFT-tel-flan overtaking TIES-tel-flan. We note that CT-then-SFT may be a better method to construct task-solving LLMs in languages with sufficient resources, for example, Bengali, Vietnamese, Indonesian, etc.
@@ 6 Understanding the Dynamics of Model Merging
Figure 4: The proportion of parameters removed in SFT-flan during sign election, relative to the total number of retained parameters of SFT-flan after trimming.
6.1 Quantification of Parameter Conflict
Revisiting the mechanism of TIES, we find that during the merging stage, parameters from one model may be discarded due to differences in the parameter signs between the two models. The discarding of model parameters usually leads to a decline in the corresponding capabilities. Considering that the CT model’s language modeling ability continuously improves with the increase of pre-training data, we suspect that it is more likely that the SFT model discards too much information during the merging process, resulting in the merged model’s inability to further enhance its task-solving ability in the target language.
To verify this hypothesis, we take Bengali, the language using the largest amount of CT corpus in our study, magnitude We explore the changes in the number of discarded parameters in the SFT model when merging it with CT models trained with different quantities of data.
@@ 6.1 Quantification of Parameter Conflict
Figure 4 illustrates the proportion of parameters that are discarded from SFT-flan during the merging stage. We can find as the LLM is pre-trained with more tokens, 4% more parameters are removed in trimmed SFT-flan. We believe that when using more tokens for CT, TIES discards a larger proportion of parameters from the SFT model, which may continuously undermine the task-solving capabilities.
@@ 6.2 Model Merging with a Slack Variable
To mitigate the information loss in SFT-flan during the merging stage, we propose TIES-SV, enhancing TIES with a Slack Variable to reduce the number of discarded parameters in the model with higher information density, i.e., SFT-flan in this situation.
@@
In conclusion, we propose a simple strategy to improve model merging. It suggests that during the process of model merging, the importance of different models’ parameters is not the same. When parameter conflicts occur, we cannot simply rely on the magnitude of the vectors to decide which model’s parameters to reserve. Instead, we should use prior information obtained through pilot studies or other means to reserve the parameters of the more important model. We hope our TIES-SV can shed light on the study of new model-merging algorithms in the future.
@@ 7 Discussions
In this section, we further explore two important questions related to model merging. First, we investigate the potential of merging more than one low-resource language into a task-solving model, which could offer a new avenue for constructing multilingual models. Second, we examine why commonly-used machine-translated data often fails in the context of low-resource languages. This failure underscores the advantage of model merging, as it does not require SFT data in the target language.
@@ 7.1 Can We Merge Multiple Languages?
Previous work Akiba et al. (2024) shows that model merging algorithms can be used to combine multiple LLMs with different capabilities, which may enhance the model to solve complex problems. We wonder whether multiple LLMs adapted to different low-resource languages can be merged with the same SFT model, to construct a a task-solving LLM supporting these languages simultaneously.
@@ 7.2 Why do Machine Translated Data Fail?
Collecting synthetic SFT data through MT is an intuitive method for constructing task-solving LLM in non-English languages Muennighoff et al. (2023); Li et al. (2023a). However, the experiment results in Table 3 show that MT-translated SFT data may not work when it comes to low-resource languages. The models trained with MT-translated data (CTSFT-X-mt) models exhibit inferior performance across all languages compared to those trained on English SFT data (CTSFT-X-flan), with a gap of -3.38$\sim$-24.61%.
The decline in model performance can be attributed to the low-quality MT results. Current open-source MT systems have limited abilities of low-resource languages. For example, we ask Uyghur and Tibetan native speakers to evaluate sampled translation results by NLLB. They find that there are often irrelevant contents and undesired code-switching in the translation results, as shown in Appendix C.2. This kind of noise not only hinders the model’s ability to learn task-solving strategies but also interferes with language modeling, which itself is inadequately learned.

### v3

@@ Abstract
Adapting large language models (LLMs) to new languages typically involves continual pre-training (CT) followed by supervised fine-tuning (SFT). However, this CT-then-SFT approach struggles with limited data in the context of low-resource languages, failing to balance language modeling and task-solving capabilities. We thus propose a new model merging solution as an alternative for low-resource languages, combining models with distinct capabilities into a single model without additional training. We use model merging to develop task-solving LLMs for low-resource languages without SFT data in the target languages. Our experiments based on Llama-2-7B demonstrate that model merging effectively endows LLMs for low-resource languages with task-solving abilities, outperforming CT-then-SFT in scenarios with extremely scarce data. Observing performance saturation in model merging with increasingly more training tokens, we further analyze the merging process and introduce a slack variable to the model merging algorithm to mitigate the loss of important parameters, thereby enhancing model performance. We hope that model merging can benefit more human languages suffering from data scarcity with its higher data efficiency.
@@ 1 Introduction
To answer these questions, we study the adaptation of Llama-2-7B Touvron et al. (2023), an English-centric LLM, into seven distinct low-resource languages. We first continually pre-train Llama-2-7B on monolingual texts in each language. Next, we explore two approaches to inject task-solving capabilities into this continually pre-trained model: (1) training the LLM with English SFT data or the data translated to the target low-resource language; (2) merging the model with an English task-solving LLM. Experiments show that model merging can effectively equip the CT models with task-solving capabilities. Notably, when pre-training corpora in the target language are extremely scarce (<10B tokens), model merging outperforms CT-then-SFT. As an LLM is continually pre-trained with more tokens in target languages, the improvements brought by model merging gradually saturate. Then, model merging can no longer significantly surpass the SFT method.
To further investigate the factors impeding the continuous improvement of model merging, we conduct a detailed analysis of the process of merging two LLMs. We find that when an English SFT model is merged with an LLM continually pre-trained with more tokens in the target language, more parameters from the SFT model are discarded during merging. The loss of these parameters may lead to a decline in task-solving capabilities, preventing the merged model from improving performance on downstream tasks. To mitigate the loss of important parameters from the SFT model, we propose a novel model merging solution with slack variables. This strategy allows for more flexible control over the merging process to retain important parameters.
@@ 3 Model Merging for Low-Resource Languages
The conventional CT-then-SFT paradigm struggles to balance language modeling and task-solving abilities in the context of low-resource languages. We propose to adopt model merging as an alternative, which can construct task-solving LLMs for low-resource languages without requiring SFT data in the target languages.
@@ 3.2 Roadmap Towards LLMs for Low-Resource Languages
Given a base model only pre-trained on an English-centric corpus, e.g., Llama-2-7B Touvron et al. (2023) in our study, we want to construct a model capable of solving tasks in a low-resource language. For those target languages, there are usually very limited pre-training texts, ranging from 1B to 20B tokens, and almost no data for supervised finetuning (SFT). In this scenario, we investigate two representative paradigms of constructing such a model: CT-then-SFT and model merging. We illustrate the roadmap in Figure 1, which demonstrates the relations among the models.
@@
The common practice is (1) first continual pre-training (CT) on the monolingual texts in the target language X to learn the language modeling and (2) then learning task-solving abilities through SFT Yong et al. (2023); Nguyen et al. (2023). This approach is referred to as CT-then-SFT. Specifically, we consider the following models:
@@
(2) CTSFT-X-mt: We translate FLAN and the training set of GSM8K into the target language X with machine translation (MT) systems [^2] We use NLLB NLLB Team et al. (2022) for translation. To enhance the model’s ability to follow English prompts, we randomly translate half of the training instances into language X, while the other half of instances remain in English., which is a common practice to obtain SFT data for non-English languages Muennighoff et al. (2023); Li et al. (2023a, b). We then finetune CT-X with the obtained SFT data.
@@
By model merging, we can integrate distinct LLMs with various capabilities into one LLM. To obtain a model capable of solving tasks in the target language X, we can merge the following two models:
CT-X: As discussed in the CT-then-SFT procedure, this model learns a certain amount of language modeling in the language X during CT. However, its task-solving ability is limited.
@@ 4 Experimental Setup
We select these languages because they are underrepresented in currently popular LLMs despite their large population (over 475M) worldwide. As shown in Table 3, the performance of Llama-2-7B in these languages is close to or even worse than random guessing [^3] The accuracy of random guessing should be 25% for Belebele, 14.29% for SIB-200, and 25% for the multiple-choice tasks in the MLiC-Eval benchmark.. Notably, the vocabulary of Llama-2 does not even contain tokens for Odia and traditional Mongolian, which indicates that the model has hardly seen these languages during pre-training. Moreover, limited resources are available for these languages on the internet. Among those languages, we can only collect fewer than 1B tokens of monolingual texts for Odia, Tibetan, Uyghur, and traditional Mongolian. The problem of data scarcity becomes more severe in terms of high-quality data for supervised fine-tuning.
@@
We translate FLAN into above mentioned languages using NLLB-200-Distilled-1.3B NLLB Team et al. (2022) [^4] NLLB is currently the open-source MT model with the most extensive language support.. Note that this model does not support traditional Mongolian and there are no open-source MT models available for this language currently. We thus adopt a roundabout way: translating the instructions into Cyrillic Mongolian, which NLLB supports, and converting them into traditional Mongolian with an open-source transliteration tool [^5] https://github.com/tugstugi/mongolian-nlp/tree/master/bichig2cyrillic.
@@
Table 2: The pre-training corpus and evaluation tasks of each language in our study. The number of tokens is obtained with the tokenizer of Llama-2.
@@
Table 3: Performance of models built through the roadmap. The best results are made bold, with the second underlined. The Task and Lang. columns denote whether the model have enhanced task-solving abilities and learned the target languages, respectively.
@@ 5.1 Effectiveness of Model Merging
In the conventional CT-then-SFT approach, LLMs often fail to acquire sufficient comprehension of the target language from a small-size CT corpus, and this ability may be further diminished by supervised fine-tuning. In contrast, model merging can preserve the language modeling acquired during CT, even with a small-size CT corpus, while incorporating task-solving capabilities by resolving parameter conflicts carefully.
@@ 5.2 Performance Plateau of Model Merging
Although model merging is shown to be more effective than CT-then-SFT in almost all languages of our study, this is not the case for Bengali, which has the largest CT corpus in our experiments: CTSFT-ben-flan outperforms TIES-ben-flan. We guess that the applicability of model merging may be related to the amount of corpus used in CT. Thus, we examine the performance changes of the merged models and the CT-then-SFT models under different amounts of CT tokens.
We collect the intermediate checkpoints of CT-ben and CT-tel, which are the two languages with the largest sizes of pre-training corpora. Then, we derive CTSFT-X-flan and TIES-X-flan models based on these checkpoints.
Figure 2 illustrates the performance of CT-only (CT-X), CT-then-SFT (CTSFT-X-flan), and model merging (TIES-X-flan) based on every checkpoint. For Bengali and Telugu, TIES-X-flan outperforms CT-X in each checkpoint, indicating model merging can robustly enhance LLM’s task-solving ability, while CT-then-SFT shows greater variability. And in both languages, for all checkpoints where the amount of pre-trained tokens is less than 10.4B tokens, TIES-X-flan can achieve better results than CTSFT-X-flan.
Model merging can integrate language modeling and task-solving capabilities more effectively than CT-then-SFT, in scenarios with limited language resources ($<$10B tokens in our experiments). Four out of seven languages in our study have pre-training corpora with fewer than 10B tokens. This data scarcity is intrinsic for low-resource languages. For instance, 81% (135 out of 166) of the languages in the multilingual corpus CulturaX Nguyen et al. (2024) have fewer than 10B tokens. Consequently, our findings on the effectiveness of model merging are broadly applicable and have the potential to benefit a wide range of human languages.
As the amount of CT tokens increases, the CTSFT-X-flan models show more rapid improvement in task-solving capabilities compared to TIE-X-flan. Specifically, when pre-trained with more than 14.4B tokens, CTSFT-ben-flan demonstrates better performance over TIES-ben-flan. Similarly, in Telugu, the performance gap between the two models slightly diminishes with additional CT tokens. However, due to the smaller size of the Telugu corpus compared to Bengali’s, we have not observed CTSFT-tel-flan overtaking TIES-tel-flan in our experiments. We note that CT-then-SFT may be a better method to construct task-solving LLMs in languages with sufficient resources, for example, Bengali, Vietnamese, Indonesian, etc.
@@ 6 Understanding the Dynamics of Model Merging
Figure 4: The proportion of parameters discarded in SFT-flan during sign election, relative to the total number of retained parameters of SFT-flan after trimming.
6.1 Quantifying Parameter Conflicts
Revisiting the mechanism of TIES, we find that during the merging stage, parameters from one model may be discarded due to differences in the parameter signs between the two models. As we continually pre-train the LLM with more tokens in language X, the parameters of CT-X changes more significantly. Since the magnitude of CT-X’s task vector becomes larger, more parameters of SFT-flan would be discarded during the process of electing signs.Figure 3 illustrates the changes of discarded parameters in SFT-flan.
The discarding of model parameters usually leads to a decline in the corresponding capabilities. Regarding that the CT model’s language modeling ability in language X continuously improves with the increase of pre-training data, we suspect that it is more likely that the SFT model is forced to discard too much information during the merging process, resulting in the merged model’s inability to further enhance its task-solving ability in the target language.
To verify this hypothesis, we take Bengali, the language with the largest amount of CT corpus in our study, as an example. We track the changes in the number of discarded parameters in the SFT model when merging it with CT models trained with different amounts of data.
@@ 6.1 Quantifying Parameter Conflicts
Figure 4 illustrates the proportion of parameters that are discarded from SFT-flan during the merging stage. We can find as the LLM is pre-trained with more tokens, 4% more parameters are removed in trimmed SFT-flan. We believe that when using more tokens for CT, TIES discards a larger proportion of parameters from the SFT model, which may continuously undermine the merged model’s task-solving capabilities .
@@ 6.2 Model Merging with a Slack Variable
To mitigate the information loss in SFT-flan during the merging stage, we design a new approach, TIES-SV, enhancing TIES with a Slack Variable to carefully reduce the number of discarded parameters in the model with higher information density, i.e., SFT-flan in this situation.
@@
It is surprising that this simple and intuitive strategy can bring such an improvement for the merged model, suggesting that different models’ parameters may have different importance during the process of model merging. When parameter conflicts occur, we cannot simply rely on the magnitude of the vectors to decide which models’ parameters should be reserve. Instead, one should use prior information obtained through pilot studies or other means to reserve the parameters of the more important model. We hope our TIES-SV can shed light on the study of new model-merging algorithms in the future.
@@ 7 Discussions
In this section, we further explore two important questions related to model merging. First, we investigate the potential of merging more than one low-resource language into a task-solving model, which could offer a new avenue for constructing multilingual models. Second, we examine why commonly-used machine-translated SFT data often fails in the context of constructing task-solving models in low-resource languages. This failure underscores the advantage of model merging, as it does not require SFT data in the target language.
@@ 7.1 Can We Merge Multiple Languages?
Previous work Akiba et al. (2024) shows that model merging algorithms can be used to combine multiple LLMs with different capabilities, which may enhance the model to solve complex problems. We wonder whether multiple LLMs adapted to different low-resource languages can be merged with the same SFT model, to construct a task-solving LLM supporting these languages simultaneously.
@@ 7.2 Why do Machine Translated Data Fail?
Collecting synthetic SFT data through MT is an intuitive method for constructing task-solving LLMs in non-English languages Muennighoff et al. (2023); Li et al. (2023a). However, the experimental results in Table 3 show that MT-translated SFT data may not work when it comes to low-resource languages. The models trained with MT-translated data (CTSFT-X-mt) exhibit inferior performance across all languages compared to those trained on English SFT data (CTSFT-X-flan), with a gap of -3.38$\sim$-24.61%.
The decline in model performance can be attributed to the low-quality MT results. Current open-source MT systems have limited abilities for low-resource languages. For example, we ask Uyghur and Tibetan native speakers to evaluate sampled translation results by NLLB. They find that there are often irrelevant contents and undesired code-switching in the translation results, as shown in Appendix C.2. This kind of noise not only hinders the model’s ability to learn task-solving strategies but also interferes with language modeling, which itself is inadequately learned.
@@
8 Case Study
We further investigate how model merging can improve the model performance on the tasks in low-resource languages. Here, we focus on two types of errors in low-resource language models: unparsable outputs during generation, and misunderstanding of character overlap in MRC. In this section, we take traditional Mongolian as an example, which the BASE model has hardly seen during pre-training.
8.1 Unparsable Output
[…]
We find CTSFT-mvf-flan produces meaningless bytes in 74 out of 200 instances, whereas TIES-mvf-flan outputs unparsable bytes in only 42 instances. Compared to the CT-then-SFT models, the merged models have better abilities to generate texts in low-resource languages.
8.2 Misunderstanding of Character Overlap
We also present a case study of the machine reading comprehension task. Here is the English translation of an instance in the MRC task of MLiC-Eval:
Context:
Man: Is the train station far from here?
Woman: It only take 20 minutes by taxi.
Question: What does the woman want to express?
Choices: A. Don’t worry. B. It is not too far.
[…]
The CT-then-SFT model demonstrates worse reasoning capabilities and tends to predict based on character overlap. We guess that fine-tuning on English datasets might primarily obtain the capabilities to handle English tasks but fail to transfer them to the low-resource languages.